Observing Log for 2007 Jan 24/25

Observers: Rick Pogge
Inst. Support: Olga Kuhn
TO: Aaron Ceranski

John Hill is also on-site through Friday.

By 11am, the dewar has leveled off at T=135K and P=7e-6mbar, so we're back to where we were before the warm up.

UTC 2007 Jan 24, 2300 - Clear and beautiful out, but the wind has been picking up. We will probably have to contend with wind and blowing snow tonight.

All times below are UTC 2007 Jan 25

0045 - Bias images (late start due to telescope work) - running 25bias.ob. Lovely sunset, beautifullly clear and photometric, but wind needs to be watched. Moon is 6 days old, sets at 0019 MST. skipping the dark image because of the late start. Because of the wind threat, we're running with the vent doors in the enclosure closed.

Data not uploading to the Repository, getting lots of "upload command failed with retcode:6" warnings in the Log Analyzer. Alex was called, eh could not reach the archive machine so we rebooted it at his instructions. Only lbcb.20070125.004730.fits made it to the archive machine before it hung.

Had to omit doing sky flats, as it was too late once we got the archive machine back online. The bias frames we took are only very slowly working their way onto the respository disk, too slow to even try keeping up with the twilight sky for flats.

0124 - at twilight std star field L95, doing initial active optics adjustment...

0140 - ready to go, L95 U @ 0140, BV @ 0144, and gr @ 0148 - seeing not too hot, 1.9 arcsec at V. Mirror is within 0.8°C

0152 - Std field PG0231, ao tweak then U @ 0203, BV @ 0206, gr @ 0212- seeing variable, 1.5-1.9 arcsec, wind up and variable, but not too fast.

0218 - Std field SA92, ao tweak then U @ 0223, BV @ 0227,gr @ 0231 - seeing creeping down to 1.5-1.6 arcsec at B

0240 - Ryan 5-deg field epoch 1 (seeing 1.6-2 arcsec), ao tweak, then go @ 0248 (note: time required for ao setup is being increased by the long time, 30-35sec, it takes for the small focus image to reach the Repository disk). Seeing on first image is ~1.5 arcsec at V, so improving.

Real paucity of early targets that can tolerate 1.5-arcsec seeing until 2100 MST, now 2000 MST. I'm going to shoot a standard star field (SA98_BV, the mistake from last night) to check the focus and seeing

0317 - SA98 seeing check and photometric standard - mirror didn't move on first try. 2nd try did it. lbcfpia barfed on one image with message
  Searching for pupils...
  Background level: 1546.58
  Computing aberrations using 17 pupils
  % MOMENT: X array must contain 2 OR more elements.
  % Error occurred at: MOMENT             99
  ...
a subsequent image did just fine.

0330 - SA98 BV shot, useful as standard and a way to measure seeing while we have a lull in the schedule. B Seeing is about 2 arcsec at 1.45 airmasses.

0343 - Ryan 5-deg field epoch 2 at 1h after the first, seeing 2" at last measure, ao tweak up first, then off @ 0349, seeing 1.6-1.8 arcsec @ V.

0405 - Going to Sigma Ori Field 1 PA=180, new OB in response to last night's data. AO tweak up, then setting up for first exposure at 0413. Will pause at the end of the first 200s image in the sequence to check seeing and scattered light from sigma Ori proper. Then decide whether to proceed. Program requires <1.5-arcsec seeing.

0419 - seeing is 1.8 arcsec at U - proceeding anyway starting at 0425 after inspecting image with OB 1 of 2. First image in sequence had 2" seeing at U. Guider is having problems locking. Not good, but stars are nice and round, so it doesn't seem to be hurting in terms of PSF shape, but seeing is 1.9-2" on the images.

0449 - Not much else to do with the seeing like this, so taking OB2 just see what that does. At least we've been getting guide stars the last couple of images so things are improving. Seeing is quite variable, from 1.4-arcsec in V to 1.8-arcsec in U. Proceeding anyway with the PI's blessings...

0529 - tweaking up primary, then executing OB3 and OB4 for Sigma Ori Field 1. Seeing still all over the place, and no 1.5-arcsec tolerant programs up yet.

0535 - executing Sigma Ori Field 1 OB3 (PA=90), U seeing still pretty awful (1.8-2 arcsec), and lack of guider lock on occasion is starting to spoil, the images (e.g., a little elongation on lbcb.20070125.053728). Not the best conditions, if still apparently photometric in terms of transparency.

0558 - executing Sigma Ori Field 1 OB4 (PA=90). Image lbcb.20070125.055444 shows jump, images doubled, guiding on really fuzzy stars. Seeing is degrading further.

0604 - humidity starting to rise now70-74%, not seeing anything blowing in a flashlight beam, operatings going up to check windspeed in the chamber. Image 055921 doesn't look so good either, and the first V image of the set has 3.2 arcsec seeing, second is 1.8 - highly variable. Stars look twinkly to the eye. Wind speed in chamber is not bad, so we're not getting obvious wind loading on the scope.

Not going to be the best data.

0632 - Ryan 5-deg field epoch 3 to finish that 2-arcsec tolerant program. Again a little later than the 1-hour interval, can't be helped. Doing mirror ao tweak then proceeding at 0638.

0654 - Midnight standards - SA98, tweaking M1 then U @ 0657, BV @ 0701, gr @ 0706 - seeing ~1.8 at B.

0711 - Standard field NGC 2419, tweaking M1 @ 0713, ... Ran into problem with lbcfpia when the LBC returned the filter ID as "B-BESSEL(?)". Trying to modify the FITS header with hedit didn't work because of how the multiext header is structured. We then hit upon the work around of loading Vfastextra.ob followed by Bfastextra.ob to reset the filter wheel. Then it worked. Meanwhile, John has modified lbcfpia to be tolerant of the (?) suffix in filter names.

0731 - wind gusting above 40mph in chamber at front of doors, so time to close for now (we heard a clunk or two in the control room).

1030 - wind situation is not improving, so we are going to call it a night.

Night Summary

At sunset the sky was a deep velvet and about as nice as I've seen anywhere. The big problem for the night was the wind, which finally rose above the 20m/s threshold and shut us down about a half hour after midnight. Seeing was awful, never better than 1.2 arcsec, and that only episodically, most of the night it hovered just below 2 arcsec. We did what we could with the programs that could tolerate bad seeing before the wind got too strong.

  • Problems with the archive computer going offline meant we could not do sky flats. After rebooting, it took many minutes before the archive machine started transferring data to the /Repository disk. The pace was so slow it would have been hopeless trying to keep up with sky flats.

  • Got photometric calibration fields since the sky was so good at sunset, but clouds started coming in at dawn.

  • Finished Ryan 5-deg field at the start of the night, which I believe finished a complete 0,5,10,15-degree set, 3 epochs each.

  • With seeing bad, we shot the 180-degree rotated Sigma Ori field 1 OBs, since the PIs indicated they could tolerate 2-arcsec seeing (which is what they got at U, better at V, sometimes). Seeing degraded and guiding became more problematic later in the sequence, so not all of the data are good. the 180-degree position and longer exposure times were used, and may help them salvage their science given the scattered light from the 3.5mag star nearby.

  • Winds went over threshold before we could get to the M81 field, which was going to be in the optimal location after finishing the midnight standard star fields. We were luckier than MMT, who was reporting wind speeds >100mph last night.

  • Had we been able to observe, our targets were the Kochanek M81 program, the Milne supernovae, and the Fan N5907 programs, all of which had a modest tolerance for poor seeing (though Milne SN2006E was a stretch). Kochanek and Milne are short enough (15min) it would be nice if they could get squeezed in during the following engineering nights with LBC, especially the Kochanek program as this gives a crucial epoch to know if the rest of the program is worth doing Feb-Apr.

  • Once the wind got over threshold, it stayed there until we finally gave up. Morning dawned with clouds and wind.

Weather has been our nemesis this run. I wish the following SDT runs better fortune.

The main instrumental problem has been the slowness of the data handling system (getting data from the instrument WinXP computer to the archive is fast, but from the archive to the Repository where we can look at it is very much slower). This cut into efficiency on things like mirror AO adjustments and sky flats. We send a report to IssueTrak on this (#452). We are experiencing 20-40 seconds to get a 1408x2304 pixel focus image to the Repository from the archive machine, compared to WinXP-to-archive rates of under a second or so from the instrument logs.

-- RichardPogge - 25 Jan 2007
Topic revision: r10 - 05 Aug 2013, KelleeSummers
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback